# Interpretable global minima of deep ReLU neural networks on sequentially separable data

### v1

@@ 4. Clustered data and dimensional reduction
Proof of proposition 4.1.
Let
$\theta_{*}:=\theta_{0}+\max_{j}\theta_{*,j},$ (64)
$P^{(\ell)}$ $=(W^{(\ell)}){}^{+}W^{(\ell)}(W^{(\ell-1)}){}^{+}W^{(\ell-1)}\cdots(W^{(1)}){}^{+}W^{(1)}$ (65)
$=(W^{(\ell)}){}^{+}W^{(\ell)}P^{(\ell-1)}.$
Induction base $\ell=1$. If we choose
$W_{1}=W^{(1)}:=\mathcal{I}(d_{0},d_{1})W_{\theta_{*}}R_{1},$ (66)
for $W_{\theta_{*}}$ defined as in (40), and $R_{1}\in O(d_{0},\mathbb{R})$ such that
$R_{1}f_{1}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert},$ (67)
and also
$b_{1}=b^{(1)}:=-W_{1}(\overline{x_{0,1}}+\mu_{1}f_{1}),$ (68)
for some $\mu_{1}$, then by lemma 3.4,
$\tau_{W^{(1)},b^{(1)}}(x)=\begin{cases}P^{(1)}x,\quad&x\in(\overline{x_{0,1}}+\mu_{1}f_{1})+\mathfrak{C}_{\theta_{1}}[f_{1}],\\ P^{(1)}(\overline{x_{0,1}}+\mu_{1}f_{1}),\quad&x\in(\overline{x_{0,1}}+\mu_{1}f_{1})+\mathfrak{C}_{\theta_{1}}[-f_{1}].\end{cases}$ (69)
$\bigcup_{j\neq 1}B_{4\delta}(\overline{x_{0,j}})\subset\overline{x_{0,1}}+\mathfrak{C}_{\theta_{*}}[f_{1}],$ (70)
and so $\mu_{1}<3\delta$ suffices [^5] This is not optimal. for
$\bigcup_{j\neq 1}B_{\delta}(\overline{x_{0,j}})\subset(\overline{x_{0,1}}+\mu_{1}f_{1})+\mathfrak{C}_{\theta_{*}}[f_{1}].$ (71)
$B_{\delta}(0)\subset-2\delta f+\mathfrak{C}_{\frac{\pi}{2}}[f]$ (72)
$B_{\delta}(\overline{x_{0,1}})\subset(\overline{x_{0,1}}+\mu_{1}f_{1})+\mathfrak{C}_{\theta_{*}}[-f_{1}].$ (73)
[…]
$\tau_{W^{(1)},b^{(1)}}(X_{0,j})=\begin{cases}P^{(1)}X_{0,j},\quad&j>1,\\ P^{(1)}(\overline{x_{0,1}}+\mu_{1}f_{1})u_{N_{1}}^{T},\quad&j=1.\end{cases}$ (74)
The induction step $\ell-1\to\ell$. Assume
$X_{0,j}^{(\tau,\ell-1)}=\begin{cases}P^{(\ell-1)}(\overline{x_{0,j}}+\mu_{j}f_{j})u_{N_{j}}^{T},\quad&j\leq\ell-1,\\ P^{(\ell-1)}X_{0,j},\quad&j>\ell-1,\end{cases}$ (75)
[…]
$X_{0,j}^{(\tau,\ell)}=\tau_{W^{(\ell)},b^{(\ell)}}(X_{0,j}^{(\tau,\ell-1)})=\begin{cases}P^{(\ell)}(\overline{x_{0,j}}+\mu_{j}f_{j})u_{N_{j}}^{T},\quad&j<\ell,\\ P^{(\ell)}(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell})u_{N_{\ell}}^{T},\quad&j=\ell,\\ P^{(\ell)}X_{0,j},\quad&j>\ell.\end{cases}$ (76)
Let
$W^{(\ell)}$ $:=\mathcal{I}(d_{0},d_{\ell})W_{\theta_{*}}R_{\ell},$
$b^{(\ell)}$ $:=-W^{(\ell)}\,(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell}),$ (77)
[…]
$R_{\ell}h_{\ell}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert}.$ (78)
$\tau_{W^{(\ell)},b^{(\ell)}}(x)=\begin{cases}(W^{(\ell)}){}^{+}W^{(\ell)}\,x,\quad&x\in(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell})+\mathfrak{C}_{\theta_{\ell}}[f_{\ell}],\\ (W^{(\ell)}){}^{+}W^{(\ell)}\,(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell}),\quad&x\in(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell})+\mathfrak{C}_{\theta_{\ell}}[-f_{\ell}].\end{cases}$ (79)
Just as in the base case, $\mu_{\ell}<3\delta$ implies
$\bigcup_{j\neq\ell}B_{\delta}(\overline{x_{0,j}})\subset(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell})+\mathfrak{C}_{\theta_{*}}[f_{\ell}],$ (80)
$B_{\delta}(\overline{x_{0,\ell}})\subset(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell})+\mathfrak{C}_{\theta_{*}}[-f_{\ell}].$ (81)
For $j<\ell$, we have from (75) that
$\mathcal{X}_{0,j}^{(\tau,\ell-1)}=P^{(\ell-1)}(\overline{x_{0,j}}+\mu_{j}f_{j}),$ (82)
and since $W^{(\ell)}P^{(\ell-1)}=W^{(\ell)}$ by (31) in lemma 3.3, we can use (43) to get
$\tau_{W^{(\ell)},b^{(\ell)}}(P^{(\ell-1)}(\overline{x_{0,j}}+\mu_{j}f_{j}))=\tau_{W^{(\ell)},b^{(\ell)}}(\overline{x_{0,j}}+\mu_{j}f_{j}).$ (83)
We claim that
$(\overline{x_{0,j}}+\mu_{j}f_{j})\in(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell})+\mathfrak{C}_{\theta_{*}}[f_{\ell}].$ (84)
Indeed, note that for all $j^{\prime}=1,\cdots,Q$,
$\overline{x}=\overline{x_{0,j^{\prime}}}+\left\lvert{\overline{x_{0,j^{\prime}}}-\overline{x}}\right\rvert f_{j^{\prime}},$ (85)
so it is clear that
$\overline{x}\in\overline{x_{0,\ell}}+\mathfrak{C}_{\theta_{*}}[f_{\ell}],$ (86)
$\overline{x}\in(\overline{x_{0,\ell}}+\mu f_{\ell})+\mathfrak{C}_{\theta_{*}}[f_{\ell}].$ (87)
By assumption (55),
$\delta<\frac{1}{4}\min_{j}\left\lvert{\overline{x_{0,j}}-\overline{x}}\right\rvert,$ (88)
so in particular
$\mu_{\ell}<3\delta<\left\lvert{\overline{x_{0,\ell}}-\overline{x}}\right\rvert.$ (89)
Again by (55) and (85),
$\mu_{j}<3\delta<\left\lvert{\overline{x_{0,j}}-\overline{x}}\right\rvert$ (90)
and $(\overline{x_{0,j}}+\mu_{j}f_{j})$ is in the line segment between $\overline{x_{0,j}}$ and $\overline{x}$, which is entirely contained in the forward cone by (80), (87) and the convexity of the cone. Having proved (84),
$=(W^{(\ell)}){}^{+}W^{(\ell)}X_{0,j}^{(\tau,\ell-1)}$
$=(W^{(\ell)}){}^{+}W^{(\ell)}P^{(\ell-1)}(\overline{x_{0,j}}+\mu_{j}f_{j})u_{N_{j}}^{T}$ (91)
$=P^{(\ell)}(\overline{x_{0,j}}+\mu_{j}f_{j})u_{N_{j}}^{T}.$
For $j=\ell$, we have assumed that
$\mathcal{X}_{0,\ell}^{(\tau,\ell-1)}=P^{(\ell-1)}\mathcal{X}_{0,\ell},$ (92)
and once again by (31),
$\tau_{W^{(\ell)},b^{(\ell)}}\left(X_{0,\ell}^{(\tau,\ell-1)}\right)=\tau_{W^{(\ell)},b^{(\ell)}}(X_{0,\ell})$ (93)
and we can use (81) in
$=(W^{(\ell)}){}^{+}W^{(\ell)}(\overline{x_{0,\ell}}+\mu_{\ell}f_{\ell})u_{N_{\ell}}^{T}$ (94)
where ()(*)( ∗ ) follows from lemma 3.3.
Similarly, for $j>\ell$,
$\mathcal{X}_{0,j}^{(\tau,\ell-1)}=P^{(\ell-1)}\mathcal{X}_{0,j},$ (95)
and we can use (31) and (80) to compute
$=(W^{(\ell)}){}^{+}W^{(\ell)}X_{0,j}^{(\tau,\ell-1)}$
$=(W^{(\ell)}){}^{+}W^{(\ell)}P^{(\ell-1)}X_{0,j}$ (96)
$=P^{(\ell)}X_{0,j},$
and we have proved (76).
Conclusion of proof. After $L-1=Q$ steps,
$X_{0,j}^{(\tau,Q)}=\overline{X_{0,j}^{(\tau,Q)}}=P^{(Q)}(\overline{x_{0,j}}-\mu_{j}f_{j})u_{N_{j}}^{T},$ (97)
so $\Delta X_{0,j}^{(\tau,Q)}=0.$ In the last layer, we have from (24)
$X^{(L)}=W^{(Q+1)}\overline{X_{0}^{(\tau,Q)}}+B^{(Q+1)}.$ (98)
$W^{(Q+1)}$ $=Y((\overline{X_{0}^{red}})^{(\tau,Q)}){}^{+},$
$b^{(Q+1)}$ $=0,$ (99)
solve
$W^{(Q+1)}\overline{X_{0}^{(\tau,Q)}}+B^{(Q+1)}=Y^{ext},$ (100)
so that the cumulative parameters (4) and (100) define a neural network with weights and biases $(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}$ such that
$X^{(L)}=Y^{ext}$ (101)
and
$\mathcal{C}[(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}]=0.$ (102)
Next, we need $P^{(Q)}$ to be suitably rank preserving. This can be done by changing the matrices $\tilde{R}$ and $R_{Q}$ in the definitions of $W_{\theta_{*}}$ and $W^{(Q)}$, respectively, as we explain now. From (30) in lemma 3.3,
$P^{(Q)}$ $=(W^{(Q)}){}^{+}W^{(Q)}$
$=(W^{(Q)})^{T}\left(W^{(Q)}(W^{(Q)})^{T}\right)^{-1}W^{(Q)}$ (103)
$(W_{\theta_{*}}R_{Q})^{T}\mathcal{I}(d_{0},d_{Q})^{T}(W^{(Q)}W^{(Q)\,T})^{-1}$ (104)
does not change the rank of any matrix on which it acts.
We now turn to the action of $\mathcal{I}(d_{0},d_{Q})W_{\theta_{*}}R_{Q}$. As we are working in barycentric coordinates (63), we only care about the action of $P^{(Q)}$ on the first $Q$ coordinates of any vector, and so restrict our attention to the first $Q$ columns of $\mathcal{I}(d_{0},d_{Q})W_{\theta_{*}}R_{Q}$. Moreover, $\mathcal{I}(d_{0},d_{Q})$ acts on $W_{\theta_{*}}R_{Q}$ by cutting off all but the first $d_{Q}\geq Q$ rows. Altogether, it is sufficient that the first $Q\times Q$ block of $W_{\theta_{*}}R_{Q}$ be invertible. Using the Laplace cofactor expansion of the determinant, it can be shown that there exists a permutation of the rows of $W_{\theta_{*}}R_{Q}$ that achieves this.
Let $P$ be the elementary matrix that realizes this permutation, and note that $P\in O(d_{0},\mathbb{R})$. Then letting $\lambda:=\lambda(\theta_{*},\theta_{d_{0}})$ in the following for notational convenience, and expanding the definition of $W_{\theta_{*}}$ from (40),
$=P\tilde{R}\operatorname{diag}(1,\lambda,\cdots,\lambda)\tilde{R}^{T}(P^{T}P)R_{Q}$ (105)
$\tilde{R}e^{d_{0}}_{1}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert}=R_{Q}f_{Q}.$ (106)
Clearly, $Pu_{d_{0}}=u_{d_{0}}$, and so we can pick $\tilde{R}$ and $R_{Q}$ satisfying (106) and such that the block $(W_{\theta_{*}}R_{Q})_{Q\times Q}$ is invertible, and $P^{(Q)}$ restricted to $\operatorname{span}\left\{\overline{x_{0,j}}\right\}_{j=1}^{Q}$ is injective.
@@ 5. Hyperplanes and sequentially separable data
$H(p,\nu):=\left\{p+x\in\mathbb{R}^{n}:\langle x,\nu\rangle=0\right\}.$ (107)
For data that is sequentially linearly separable, we have the corresponding hyperplanes $H_{j}=H(p_{j},h_{j})$ for some unit vectors $h_{j}$, $j=1,\cdots,Q$. We will always consider data sets to be finite, and so the distance from any such hyperplane to the each of the two sets it is separating is always positive. Moreover, a pair $(p,\nu)\subset\mathbb{R}^{n}\times\mathbb{R}^{n}$, along with an angle $\theta<\pi$, define a pair of cones $p+\mathfrak{C}_{\theta}[\nu]$ and $p+\mathfrak{C}_{\theta}[-\nu]$ as in (35), and so to each hyperplane $H(p,\nu)$ is associated a family of cones.
@@
$\min_{(W_{i},b_{i})_{i=1}^{L}}\mathcal{C}[(W_{i},b_{i})_{i=1}^{L}]=0,$ (108)
with weights and biases given recursively by (21), (22) for the following cumulative parameters: for $\ell=1,\cdots,Q$,
@@
$\bigcup_{j>1}\mathcal{X}_{0,j}$ $\subset p_{1}+\mathfrak{C}_{\theta_{1}}[h_{1}].$ (113)
@@
$W_{1}=W^{(1)}:=W_{\theta_{1}}R_{1},$ (114)
for $W_{\theta_{1}}$ defined as in (40), and $R_{1}\in O(d_{0},\mathbb{R})$ such that
$R_{1}h_{1}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert},$ (115)
@@
$\tau_{W^{(1)},b^{(1)}}(x)=\begin{cases}x,\quad&x\in p_{1}+\mathfrak{C}_{\theta_{1}}[h_{1}],\\ p_{1},\quad&x\in p_{1}+\mathfrak{C}_{\theta_{1}}[-h_{1}],\end{cases}$ (117)
@@
$X_{0,j}^{(\tau,1)}=\tau_{W^{(1)},b^{(1)}}(X_{0,j})=\begin{cases}p_{1}u_{N_{1}}^{T},\quad&j=1,\\ X_{0,j},\quad&j>1,\end{cases}$ (118)
@@
$X_{0,j}^{(\tau,\,\ell-1)}=\begin{cases}p_{j}u_{N_{j}}^{T},\quad&j\leq\ell-1,\\ X_{0,j},\quad&j>\ell-1,\end{cases}$ (119)
where $p_{j}$ are the points given by (111). We will construct cumulative parameters $W^{(\ell)},b^{(\ell)}$ such that
$X_{0,j}^{(\tau,\ell)}=\tau_{W^{(\ell)},b^{(\ell)}}(X_{0,j}^{(\tau,\ell-1)})=\begin{cases}p_{j}u_{N_{j}}^{T}=X_{0,j}^{(\tau,\ell-1)},\quad&j<\ell,\\ p_{\ell}u_{N_{\ell}}^{T},\quad&j=\ell,\\ X_{0,j}=X_{0,j}^{(\tau,\ell-1)},\quad&j>\ell.\end{cases}$ (120)
By sequential linear separability of the data, there exists a hyperplane $H_{\ell}=H(p_{\ell},h_{\ell})$ in the collection given by (111) separating $\mathcal{X}_{0,\ell}$ and $\bigcup_{j<\ell}\{p_{j}\}\cup\bigcup_{j>\ell}\mathcal{X}_{0,j}$. We can use the induction hypothesis (119) to rewrite
$\mathcal{X}_{0,\ell}=\mathcal{X}_{0,\ell}^{(\tau,\ell-1)}\,\text{and }\,\bigcup_{j<\ell}\{p_{j}\}\cup\bigcup_{j>\ell}\mathcal{X}_{0,j}=\bigcup_{j\neq\ell}\mathcal{X}_{0,j}^{(\tau,\ell-1)},$ (121)
@@
$\bigcup_{j\neq\ell}\mathcal{X}_{0,j}^{(\tau,\ell-1)}\subset p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[h_{\ell}]\quad\text{ and }\quad\mathcal{X}_{0,\ell}\subset p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[-h_{\ell}].$ (122)
@@
$b^{(\ell)}:=-W^{(\ell)}\,p_{\ell},$ (123)
for $W_{\theta_{\ell}}$ defined as in (40) and $R_{\ell}\in O(n,\mathbb{R})$ such that
$R_{\ell}h_{\ell}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert},$ (124)
@@
$\tau_{W^{(\ell)},b^{(\ell)}}(x)=\begin{cases}x,\quad&x\in p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[h_{\ell}],\\ p_{\ell},\quad&x\in p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[-h_{\ell}],\end{cases}$ (125)
which, along with (122), implies (120).
@@
$\min_{W^{(Q+1)},b^{(Q+1)}}\left\lVert{W^{(Q+1)}(X_{0})^{(\tau,Q)}+B^{(Q+1)}-Y^{ext}}\right\rVert^{2}_{\mathcal{L}^{2}}.$ (127)
@@
$(\overline{X_{0}^{red}})^{(\tau,Q)}=\left[p_{1}\cdots p_{Q}\right]$ (132)
@@
$((\overline{X_{0}^{red}})^{(\tau,Q)}){}^{+}(\overline{X_{0}^{red}})^{(\tau,Q)}={\textbf{1}}_{Q\times Q},$ (133)
@@
$W^{(Q+1)}=Y((\overline{X_{0}^{red}})^{(\tau,Q)}){}^{+}$ (134)
is a solution to (131), and we have in fact produced a collection of weights and biases $(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}$ given by the cumulative parameters defined in (5) and (134) that achieves zero cost,
$\mathcal{C}[(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}]=0.$ (135)

### v2

@@ 4. Clustered data and dimensional reduction
We defer the full proof to section 6.
@@ 5. Hyperplanes and sequentially separable data
$H(p,\nu):=\left\{p+x\in\mathbb{R}^{n}:\langle x,\nu\rangle=0\right\}.$ (5.1)
For data that is sequentially linearly separable, we have the corresponding hyperplanes $H_{j}=H(p_{j},h_{j})$ for some unit vectors $h_{j}$, $j=1,\cdots,Q$. We will always consider data sets to be finite, and so the distance from any such hyperplane to the each of the two sets it is separating is always positive. Moreover, a pair $(p,\nu)\subset\mathbb{R}^{n}\times\mathbb{R}^{n}$, along with an angle $\theta<\pi$, define a pair of cones $p+\mathfrak{C}_{\theta}[\nu]$ and $p+\mathfrak{C}_{\theta}[-\nu]$ as in (3.18), and so to each hyperplane $H(p,\nu)$ is associated a family of cones.
@@
$\min_{(W_{i},b_{i})_{i=1}^{L}}\mathcal{C}[(W_{i},b_{i})_{i=1}^{L}]=0,$ (5.2)
with weights and biases given recursively by (3.4), (3.5) for the following cumulative parameters: for $\ell=1,\cdots,Q$,
@@
$\bigcup_{j>1}\mathcal{X}_{0,j}$ $\subset p_{1}+\mathfrak{C}_{\theta_{1}}[h_{1}].$ (5.7)
@@
$W_{1}=W^{(1)}:=W_{\theta_{1}}R_{1},$ (5.8)
for $W_{\theta_{1}}$ defined as in (3.23), and $R_{1}\in O(d_{0},\mathbb{R})$ such that
$R_{1}h_{1}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert},$ (5.9)
@@
$\tau_{W^{(1)},b^{(1)}}(x)=\begin{cases}x,\quad&x\in p_{1}+\mathfrak{C}_{\theta_{1}}[h_{1}],\\ p_{1},\quad&x\in p_{1}+\mathfrak{C}_{\theta_{1}}[-h_{1}],\end{cases}$ (5.11)
@@
$X_{0,j}^{(\tau,1)}=\tau_{W^{(1)},b^{(1)}}(X_{0,j})=\begin{cases}p_{1}u_{N_{1}}^{T},\quad&j=1,\\ X_{0,j},\quad&j>1,\end{cases}$ (5.12)
@@
$X_{0,j}^{(\tau,\,\ell-1)}=\begin{cases}p_{j}u_{N_{j}}^{T},\quad&j\leq\ell-1,\\ X_{0,j},\quad&j>\ell-1,\end{cases}$ (5.13)
where $p_{j}$ are the points given by (5.5). We will construct cumulative parameters $W^{(\ell)},b^{(\ell)}$ such that
$X_{0,j}^{(\tau,\ell)}=\tau_{W^{(\ell)},b^{(\ell)}}(X_{0,j}^{(\tau,\ell-1)})=\begin{cases}p_{j}u_{N_{j}}^{T}=X_{0,j}^{(\tau,\ell-1)},\quad&j<\ell,\\ p_{\ell}u_{N_{\ell}}^{T},\quad&j=\ell,\\ X_{0,j}=X_{0,j}^{(\tau,\ell-1)},\quad&j>\ell.\end{cases}$ (5.14)
By sequential linear separability of the data, there exists a hyperplane $H_{\ell}=H(p_{\ell},h_{\ell})$ in the collection given by (5.5) separating $\mathcal{X}_{0,\ell}$ and $\bigcup_{j<\ell}\{p_{j}\}\cup\bigcup_{j>\ell}\mathcal{X}_{0,j}$. We can use the induction hypothesis (5.13) to rewrite
$\mathcal{X}_{0,\ell}=\mathcal{X}_{0,\ell}^{(\tau,\ell-1)}\,\text{and }\,\bigcup_{j<\ell}\{p_{j}\}\cup\bigcup_{j>\ell}\mathcal{X}_{0,j}=\bigcup_{j\neq\ell}\mathcal{X}_{0,j}^{(\tau,\ell-1)},$ (5.15)
@@
$\bigcup_{j\neq\ell}\mathcal{X}_{0,j}^{(\tau,\ell-1)}\subset p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[h_{\ell}]\quad\text{ and }\quad\mathcal{X}_{0,\ell}\subset p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[-h_{\ell}].$ (5.16)
@@
$b^{(\ell)}:=-W^{(\ell)}\,p_{\ell},$ (5.17)
for $W_{\theta_{\ell}}$ defined as in (3.23) and $R_{\ell}\in O(n,\mathbb{R})$ such that
$R_{\ell}h_{\ell}=\frac{u_{d_{0}}}{\left\lvert{u_{d_{0}}}\right\rvert},$ (5.18)
@@
$\tau_{W^{(\ell)},b^{(\ell)}}(x)=\begin{cases}x,\quad&x\in p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[h_{\ell}],\\ p_{\ell},\quad&x\in p_{\ell}+\mathfrak{C}_{\theta_{\ell}}[-h_{\ell}],\end{cases}$ (5.19)
which, along with (5.16), implies (5.14).
@@
$\min_{W^{(Q+1)},b^{(Q+1)}}\left\lVert{W^{(Q+1)}(X_{0})^{(\tau,Q)}+B^{(Q+1)}-Y^{ext}}\right\rVert^{2}_{\mathcal{L}^{2}}.$ (5.21)
@@
$(\overline{X_{0}^{red}})^{(\tau,Q)}=\left[p_{1}\cdots p_{Q}\right]$ (5.26)
@@
$((\overline{X_{0}^{red}})^{(\tau,Q)}){}^{+}(\overline{X_{0}^{red}})^{(\tau,Q)}={\textbf{1}}_{Q\times Q},$ (5.27)
@@
$W^{(Q+1)}=Y((\overline{X_{0}^{red}})^{(\tau,Q)}){}^{+}$ (5.28)
is a solution to (5.25), and we have in fact produced a collection of weights and biases $(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}$ given by the cumulative parameters defined in (5) and (5.28) that achieves zero cost,
$\mathcal{C}[(W_{i}^{**},b_{i}^{**})_{i=1}^{Q+1}]=0.$ (5.29)